Even if both the partial derivatives of R → R
2
f : exist at a point )b,a( , the
function f may still not be continuous at )b,a( .
1
Expert's answer
2020-06-14T18:28:19-0400

#Q121564

Now the following example satisfied the given condition

Exp:

f(x,y)= xy /(x2+y2)

Now let us find partial derivative of f(x,y) with respect to x at (0,0)

fx = y / (x2 +y2)

So, fx(0,0) = 0

Now let us find partial derivative of f(x,y) with respect to y at (0,0)

fy = x / (x2+y2)

So, fy(0,0)=0

Therefore , both partial derivative of f(x,y) exist at (0,0)

Now let us check continuity of f(x,y) at (0,0)

Put y=mx into f(x,y)

So, f(x,y)= mx2/x2.(1+m2) = m/(1+m2)

Hence , limt of f(x,y) as (x,y) "\to" (0,0)

Depends upon the value of m

So, limit doesn't exist

Hence , f(x,y) is not continuous at (0,0)

Ans. f(x,y)= xy /(x2+y2) satisfied given condition.